# Finding separatrices of dynamical flows with Deep Koopman Eigenfunctions

### v1

@@ 4 Discussion
Prior efforts in reverse-engineering RNNs relied heavily on locating fixed points and linearizing dynamics locally [5, 6, 10, 12, 7, 8, 9, 11]. While powerful, these methods cannot directly capture global structures or predict system responses to large perturbations that cross basin boundaries. By directly approximating scalar-valued KEFs that vanish precisely on separatrices, our method complements and extends existing local linearization approaches. Practitioners can use our KEFs alongside fixed-point analysis to achieve a comprehensive understanding of the dynamical system’s landscape.
@@
Our work also advances the application of Koopman operator theory to dynamical systems. Previous studies primarily utilized Koopman eigenfunctions to predict or control dynamics within a single basin of attraction [23, 32, 33, 34, 35]. Likewise, methods comparing dynamical systems to one another use the dynamic mode decomposition which does not always discern between different basins [14, 15, 36]. Such studies usually involve KEFs associated with negative eigenvalues ($\lambda$<0), which exhibit opposite behavior to ours: they explode at separatrices and approach zero at attractors. In contrast, we specifically targeted eigenfunctions associated with positive eigenvalues, ensuring their zeros correspond exactly to separatrices.
@@
To help practitioners use our method, we highlight inherent challenges, such as degeneracy in the Koopman PDE, where multiple solutions exist. To overcome these, we introduced regularization strategies, notably a balance term ensuring eigenfunctions change sign across different basins. Additionally, ensemble training and targeted sampling methods were used to resolve directional ambiguities and ensure comprehensive coverage of separatrix geometry. These mitigations join existing work on KEF approximation [25, 24, 37] and enabled reliable identification of separatrices in diverse and high-dimensional systems.
@@
Beyond the conceptual aspect of using positive KEFs to extract separatrices, and the qualitative manners in which this procedure can go wrong, there is also a computational aspect. While we did not elaborate on this here, it is important to note that finding the KEF is an iterative procedure in phase space. In contrast, it is possible to use grid searches and bisections to simulate the ODE in many locations to search for the separatrix [29]. The computational load of the latter approach scales with time, and requires to compute the dynamics associated with the same area in phase space many times. In contrast, solving the PDE is a form of dynamic programming that can be made more efficient. Specifically for the task of locating separatrices, critical slowing down can make solving the ODE computationally heavy.
@@
While we demonstrated the applicability of our method to diverse scenarios, we do not provide theoretical guarantees linking the accuracy of the KEF approximation and that of the separatrix location. Furthermore, like techniques for finding fixed points [5, 38], our method requires knowing the dynamics in the entire phase space. In the future, extending this to trajectory-based methods [39, 40, 14] can facilitate the application of the method in neuroscience settings.
@@ Appendix A Analytical KEF derivation in 1D bistable system
[…]
@@ Appendix C Relation of our definition to the Koopman Operator
$(\mathcal{K}_{\tau}g)(\boldsymbol{x}(t))=g(F_{\tau}(\boldsymbol{x}(t)))=g(\boldsymbol{x}(t+\tau)).$ (44)
@@
$\mathcal{K}g:=\lim_{\tau\to 0}\frac{\mathcal{K}_{\tau}g-g}{\tau}=\lim_{\tau\to 0}\frac{g(F_{\tau}(\boldsymbol{x}))-g(\boldsymbol{x})}{\tau}.$ (45)
@@
$\mathcal{K}g(\boldsymbol{x}(t))=\lim_{\tau\to 0}\frac{g(\boldsymbol{x}(t+\tau))-g(\boldsymbol{x}(t))}{\tau}=\frac{d}{dt}g(\boldsymbol{x}(t))=\nabla g\cdot\boldsymbol{\dot{x}}(t)=\nabla g\cdot f\big{(}\boldsymbol{x}(t)\big{)}.$ (46)
@@ Appendix E Curve-based validation approach
The tangent vectors $m_{x}$ and $m_{y}$ are initialized as $y-x$ and perturbed with Gaussian noise:
@@ Appendix F Neural network architectures
In most of the demonstrations we use a ResNet architecture [41] with a tanh activation function.
Let the input to the network be $\mathbf{x}_{\text{in}}\in\mathbb{R}^{d_{\text{in}}}$ and define the hidden layer activations $\mathbf{x}_{\text{hid}}\in\mathbb{R}^{d_{\text{hid}}}$, output dimension $\mathbf{x}_{\text{out}}\in\mathbb{R}^{d_{\text{out}}}$, and number of layers $L$.
The network receives inputs at the first layer $\mathbf{x}^{(0)}=\text{Pad}(\mathbf{x}_{\text{in}}$), where Pad appends zeroes to the input as we always choose $d_{\text{hid}}>d_{\text{in}}$. The network then transforms the input at each layer $l$,
@@
For the results in Figures 2, 4, 7, we use $L=20$, $d_{\text{hid}}=400$. $d_{\text{out}}=1$ and $d_{\text{in}}=N$ the dimension of the dynamical system. For the gLV system in Figure 5 we use $L=25$ and $d_{\text{hid}}=1000$.
@@
For the limit cycles example in Figure 6 we use a single Radial Basis Function layer [30].
@@ Appendix G Optimisation
where $j$ corresponds to the $j^{\text{th}}$ sampling distribution (see main text section 3.2). $B$ $N$-dimensional points in the state space $\mathcal{X}$ are sampled from each distribution $\boldsymbol{x}^{j}\sim p_{j}(\boldsymbol{x})$. The ratio loss is the Koopman PDE error, normalised by a sample-shuffled version:
@@
We use the Adam optimiser [31] with learning rate $10^{-4}$ and l2 normalisation $10^{-5}$. We use $B=1000$ and train for $1000$ iterations.
@@
Table 1: Algorithm details and hyperparameters for various systems. System dimensionality $N$, Koopman eigenvalue $\lambda$, balance regularisation weight $\gamma_{\text{bal}}$, batch-size $B$, training iterations $T$, learning rate $\eta$, ResNet depth $L$ and width $d_{\text{hid}}$, number of Radial Basis Functions $M$.
Dynamical System $N$ $\lambda$ $\gamma_{\text{bal}}$ $B$ $T$ $\eta$ $L$ $d_{\text{hid}}$ $M$
Bistable 1D $1$ $1$ $5\times 10^{-2}$ $1000$ $1000$ $10^{-4}$ $20$ $400$
Damped Duffing oscillator $2$ $1$ $5\times 10^{-2}$ $1000$ $1000$ $10^{-4}$ $20$ $400$
1BFF, 2D GRU $2$ $1$ $5\times 10^{-2}$ $1000$ $1000$ $10^{-4}$ $20$ $400$
2BFF, 3D GRU $3$ $0.2$ $5\times 10^{-2}$ $1000$ $1000$ $10^{-4}$ $20$ $400$
1BFF, 64D $64$ $0.1$ $5\times 10^{-2}$ $1000$ $1000$ $10^{-4}$ $20$ $400$
Two Limit Cycles $2$ $1$ 00 $1000$ $1000$ $10^{-4}$ $300$
Ecology gLV [28, 29] $11$ $0.1$ $5\times 10^{-2}$ $5000$ $5000$ $10^{-4}$ $25$ $1000$
Appendix H Computational Resources and run time

### v2

@@ 4 Discussion
Prior efforts in reverse-engineering RNNs relied heavily on locating fixed points and linearizing dynamics locally [carnevale_dynamic_2015, maheswaranathan_reverse_2019, maheswaranathan_universality_2019, finkelstein_attractor_2021, mante_context-dependent_2013, liu_encoding_2024, driscoll_flexible_2024, low_remapping_2023, chaisangmongkon_computing_2017, pagan_individual_2025, wang_flexible_2018]. While powerful, these methods cannot directly capture global structures or predict system responses to large perturbations that cross basin boundaries. By directly approximating scalar-valued KEFs that vanish precisely on separatrices, our method complements and extends existing local linearization approaches. Practitioners can use our KEFs alongside fixed-point analysis to achieve a comprehensive understanding of the dynamical system’s landscape.
While FTLE methods [haller_lagrangian_2000, tanaka_separatrices_2009, haller_lagrangian_2015] also identify separatrices as ridges of finite-time trajectory divergence, they may change sharply near separatrices while providing little gradient elsewhere. We speculate that this could limit their usefulness in high-dimensional systems, where gradient-based localization is needed. Moreover, differentiating FTLE requires differentiating through the dynamical function, which may be computationally expensive or even infeasible due to vanishing/exploding gradients. In contrast, our Koopman eigenfunction framework, by integrating globally, provides a smooth scalar field whose zero level set identifies the separatrix, enabling efficient gradient-based searches without repeated forward simulations of the target system.
@@
Our work also advances the application of Koopman operator theory to dynamical systems. Previous studies primarily utilized Koopman eigenfunctions to predict or control dynamics within a single basin of attraction [lusch_deep_2018, cohen_latent_2023, mezic_spectral_2005, mezic_koopman_2016, meng_koopman-based_2024]. Likewise, methods comparing dynamical systems to one another use the dynamic mode decomposition which does not always discern between different basins [ostrow_beyond_2023, redman_identifying_2024, mezic_comparison_2004]. Such studies usually involve KEFs associated with negative eigenvalues ($\lambda$<0), which exhibit opposite behavior to ours: they explode at separatrices and approach zero at attractors. In contrast, we specifically targeted eigenfunctions associated with positive eigenvalues, ensuring their zeros correspond exactly to separatrices.
@@
To help practitioners use our method, we highlight inherent challenges, such as degeneracy in the Koopman PDE, where multiple solutions exist. To overcome these, we introduced regularization strategies, notably a balance term ensuring eigenfunctions change sign across different basins. Additionally, ensemble training and targeted sampling methods were used to resolve directional ambiguities and ensure comprehensive coverage of separatrix geometry. These mitigations join existing work on KEF approximation [deka_koopman-based_2022, yeung_learning_2019, deka_supervised_2023] and enabled reliable identification of separatrices in diverse and high-dimensional systems.
@@
Beyond the conceptual aspect of using positive KEFs to extract separatrices, and the qualitative manners in which this procedure can go wrong, there is also a computational aspect. While we did not elaborate on this here, it is important to note that finding the KEF is an iterative procedure in phase space. In contrast, it is possible to use grid searches and bisections to simulate the ODE in many locations to search for the separatrix [jones_steady-state_2019]. The computational load of the latter approach scales with time, and requires computing the dynamics associated with the same area in phase space many times. In contrast, solving the PDE is a form of dynamic programming that can be made more efficient (see Appendix H). Specifically for the task of locating separatrices, critical slowing down can make solving the ODE computationally demanding.
@@
While we demonstrated the applicability of our method to diverse scenarios, we do not provide theoretical guarantees linking the accuracy of the KEF approximation and that of the separatrix location. Furthermore, like techniques for finding fixed points [sussillo_opening_2013, katz_using_2018], our method requires knowing the dynamics in the entire phase space. In the future, extending this to trajectory-based methods [turner_charting_2021, brennan_one_2023, ostrow_beyond_2023] can facilitate the application of the method in neuroscience settings.
An interesting extension of our work is to stochastic dynamical systems [linderman_bayesian_2017, duncker_learning_2019]. While one can approximate separatrices using only the deterministic component of the dynamics, a full treatment of stochasticity and its effect on basin boundaries remains open. Likewise, data-driven models with uncertainty in the inferred flow [hu_modeling_2024] raise questions about how this uncertainty propagates to the separatrix. Since separatrices often occupy sparsely sampled or unstable regions of phase space, this motivates active sampling—potentially through targeted optogenetic stimulations—to improve model accuracy along these boundaries [wagenmaker_active_2024].
@@
Acknowledgements
This work was supported by the Israel Science Foundation (grant No. 1442/21 to OB) and Human Frontiers Science Program (HFSP) research grant (RGP0017/2021 to OB). The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
@@ Appendix A Analytical KEF derivation in 1D bistable system
$\int\frac{1}{x-x^{3}}\,dx$ $=\int\left(\frac{1}{x}+\frac{1}{2(1-x)}-\frac{1}{2(1+x)}\right)dx$ (20)
@@ Appendix C Relation of our definition to the Koopman Operator
Remark. When studying nonlinear dynamical systems with multiple basins of attraction, as we do, the corresponding Koopman eigenfunctions are generally not square-integrable, and therefore fall outside the Hilbert space defined by the standard inner product. We are aware of this theoretical limitation and continue to employ the Koopman framework regardless. In practice, neural networks learn finite, smooth approximations to these otherwise singular structures.
@@
$(\mathcal{K}_{\tau}g)\bigl(\boldsymbol{x}(t)\bigr)=g\Bigl(F_{\tau}\bigl(\boldsymbol{x}(t)\bigr)\Bigr)=g\bigl(\boldsymbol{x}(t+\tau)\bigr).$ (41)
@@
$\mathcal{K}g:=\lim_{\tau\to 0}\frac{\mathcal{K}_{\tau}g-g}{\tau}=\lim_{\tau\to 0}\frac{g\Bigl(F_{\tau}\bigl(\boldsymbol{x}\bigr)\Bigr)-g\bigl(\boldsymbol{x}\bigr)}{\tau}.$ (42)
@@
$\mathcal{K}g\bigl(\boldsymbol{x}(t)\bigr)$ $=\lim_{\tau\to 0}\frac{g\bigl(\boldsymbol{x}(t+\tau)\bigr)-g\bigl(\boldsymbol{x}(t)\bigr)}{\tau}$ (43)
$=\frac{d}{dt}\,g\bigl(\boldsymbol{x}(t)\bigr)=\nabla g\bigl(\boldsymbol{x}(t)\bigr)\cdot\boldsymbol{\dot{x}}(t)=\nabla g\bigl(\boldsymbol{x}(t)\bigr)\cdot f\Bigl(\boldsymbol{x}(t)\Bigr).$ (44)
@@ Appendix E Curve-based validation approach
The tangent vectors $m_{x}$ and $m_{y}$ are initialized as yxy-x and perturbed with Gaussian noise:
@@ Appendix F Neural network architectures
In most of the demonstrations we use a ResNet architecture [he_deep_2016] with a tanh activation function.
Let the input to the network be $\mathbf{x}_{\text{in}}\in\mathbb{R}^{d_{\text{in}}}$ and define the hidden layer activations $\mathbf{x}_{\text{hid}}\in\mathbb{R}^{d_{\text{hid}}}$, output dimension $\mathbf{x}_{\text{out}}\in\mathbb{R}^{d_{\text{out}}}$, and number of layers $L$. In some cases Let the input to the network be $\mathbf{x}_{\text{in}}\in\mathbb{R}^{d_{\text{in}}}$ and define the hidden layer activations $\mathbf{x}_{\text{hid}}\in\mathbb{R}^{d_{\text{hid}}}$, output dimension $\mathbf{x}_{\text{out}}\in\mathbb{R}^{d_{\text{out}}}$, and number of layers $L$.
The network receives inputs at the first layer $\mathbf{x}^{(0)}=\text{Pad}(a\mathbf{x}_{\text{in}}$), where Pad appends zeroes to the input as we always choose $d_{\text{hid}}>d_{\text{in}}$, and $a\in\mathbb{R}_{+}$ is a scalar that we set to ensure elements of $\mathbf{x}^{(0)}$ are $\mathcal{O}(1)$. The network then transforms the input at each layer $l$,
@@
For the results in Figures 3, 5, 8, we use $L=20$, $d_{\text{hid}}=400$. $d_{\text{out}}=1$ and $d_{\text{in}}=N$ the dimension of the dynamical system. For the gLV system in Figure 6 we use $L=25$ and $d_{\text{hid}}=1000$.
@@
For the limit cycles example in Figure 7 we use a single Radial Basis Function layer [powell_radial_1987].
@@ Appendix G Optimisation
Our optimisation procedure described in the main text is summarised in Algorithm 1. It implements the training of the neural network Koopman eigenfunction using the ratio loss and balance regularisation, for multiple sampling distributions.
Algorithm 1 Train Koopman Eigenfunction Network
1:Sampling distributions $\{p_{j}(\boldsymbol{x})\}_{j=1}^{J}$; vector field $f$; eigenvalue $\lambda$; neural network architecture $\psi_{\theta}$; batch size $B$; iterations $T$; balance weight $\gamma_{\mathrm{bal}}$; learning rate $\eta$; small $\varepsilon=10^{-12}$
2:Initialize $\theta$
3:for $t=1\to T$ do
4:  $L_{\text{total}}\leftarrow 0$
5:  for $j=1\to J$ do
6:   $\{\boldsymbol{x}_{i}^{j},\psi_{i}^{j},\nabla\psi_{i}^{j}\}\leftarrow$ SampleAndEvaluate($p_{j},\psi_{\theta},B$)
[…]
8:   $L_{\text{bal}}^{j}\leftarrow$ ComputeBalanceLoss($\{\psi_{i}^{j}\}$)
9:   $L_{\text{total}}\leftarrow L_{\text{total}}+L_{\text{ratio}}^{j}+\gamma_{\mathrm{bal}}\,L_{\text{bal}}^{j}$
10:  end for
11:  Compute gradients of $L_{\text{total}}$ w.r.t. $\theta$
12:  Update weights $\theta$ using gradients and learning rate $\eta$
13:end for
14:return trained parameters $\theta$
15:procedure SampleAndEvaluate($p_{j},\psi_{\theta},B$)
16:  Sample $\{\boldsymbol{x}_{i}^{j}\}_{i=1}^{B}\sim p_{j}(\boldsymbol{x})$
17:  Compute $\psi_{i}^{j}\leftarrow\psi_{\theta}(\boldsymbol{x}_{i}^{j})$
18:  Compute $\nabla\psi_{i}^{j}\leftarrow\nabla_{\boldsymbol{x}}\psi_{\theta}(\boldsymbol{x}_{i}^{j})$
19:  return $\{\boldsymbol{x}_{i}^{j},\psi_{i}^{j},\nabla\psi_{i}^{j}\}$
20:end procedure
[…]
22:  Compute $\mathrm{LHS}_{i}^{j}=\nabla\psi_{i}^{j}\cdot f(\boldsymbol{x}_{i}^{j})$
23:  Compute $\mathrm{RHS}_{i}^{j}=\lambda\,\psi_{i}^{j}$
24:  Draw random permutation $\pi$ of $\{1,\dots,B\}$
25:  $n_{j}\leftarrow\sum_{i}(\mathrm{LHS}_{i}^{j}-\mathrm{RHS}_{i}^{j})^{2}$
26:  $d_{j}\leftarrow\sum_{i}(\mathrm{LHS}_{i}^{j}-\mathrm{RHS}_{\pi(i)}^{j})^{2}$
27:  return $L_{\text{ratio}}^{j}=n_{j}/(d_{j}+\varepsilon)$
28:end procedure
29:procedure ComputeBalanceLoss($\{\psi_{i}^{j}\}$)
30:  $\bar{\psi}^{j}=\frac{1}{B}\sum_{i}\psi_{i}^{j}$
31:  $v^{j}=\frac{1}{B}\sum_{i}(\psi_{i}^{j}-\bar{\psi}^{j})^{2}$
32:  return $L_{\text{bal}}^{j}=(\bar{\psi}^{j})^{2}/(v^{j}+\varepsilon)$
33:end procedure
@@
where $j$ corresponds to the $j^{\text{th}}$ sampling distribution (see main text section 3.2). $B$ NN-dimensional points in the state space $\mathcal{X}$ are sampled from each distribution $\boldsymbol{x}^{j}_{i}\sim p_{j}(\boldsymbol{x})$. The ratio loss is the Koopman PDE error, normalised by a sample-shuffled version:
@@
We use the Adam optimiser [kingma_adam_2017] with learning rate $10^{-4}$ and l2 normalisation $10^{-5}$. We use $B=1000$ and train for $1000$ iterations.
@@
Table 1: Algorithm details and hyperparameters for various systems. System dimensionality NN, Koopman eigenvalue $\lambda$, balance regularisation weight $\gamma_{\text{bal}}$, batch-size $B$, training iterations $T$, learning rate $\eta$, ResNet depth $L$ and width $d_{\text{hid}}$, number of Radial Basis Functions $M$.
Dynamical System NN $\lambda$ $\gamma_{\text{bal}}$ $B$ $T$ $\eta$ $L$ $d_{\text{hid}}$ $M$
Bistable 1D 11 11 $0.05$ $1000$ $1000$ $10^{-4}$ 2020 $400$
Damped Duffing oscillator 22 11 $0.05$ $1000$ $1000$ $10^{-4}$ 2020 $400$
1BFF, 2D GRU 22 11 $0.05$ $1000$ $1000$ $10^{-4}$ 2020 $400$
2BFF, 3D GRU 33 $0.2$ $0.05$ $1000$ $1000$ $10^{-4}$ 2020 $400$
1BFF, 64D $64$ $0.1$ $0.05$ $1000$ $1000$ $10^{-4}$ 2020 $400$
Two Limit Cycles 22 11 0 $1000$ $1000$ $10^{-4}$ $300$
Ecology gLV $11$ $0.1$ $0.05$ $5000$ $5000$ $10^{-4}$ $25$ $1000$
Data-trained RNN [finkelstein_attractor_2021] $668$ $0.02$ $0.05$ $1000$ $1000$ $10^{-4}$ 77 $1200$
G.1 Choice of Training Distributions
Choosing suitable training distributions $p_{j}(\boldsymbol{x})$ is an important step in applying our method. The distributions are selected to satisfy the following criteria:
  1. 1.
    They are approximately bisected by the separatrix, ensuring roughly equal sampling from both basins and enabling the balance loss to be satisfied.
  2. 2.
    They sample sufficiently near the attractors to capture the global bistable dynamics.
  3. 3.
    They are approximately forward-invariant, i.e., trajectories initialized from the distribution remain within its support when evolved forward in time. This avoids loss of mass due to transient amplification along unstable directions.
  4. 4.
    In systems with multiple spurious attractors, they avoid sampling from basins outside the domain of interest.
Below we list the specific distributions used for each system.
1D bistable system (Fig. 3C): $\mathcal{N}(0,1),\,\mathcal{N}(0,3)]$.
2D damped Duffing oscillator (Fig. 3F): $\mathcal{N}(\boldsymbol{0},\,\sigma_{j}^{2}\mathbf{I}_{2})$, with $\sigma_{j}=[0.1,0.5,1.0,2.0]$, where $\mathbf{I}_{2}$ is the $2\times 2$ identity matrix.
2D 1-bit flip-flop GRU (Fig. 3I): $\mathcal{N}(\boldsymbol{0},\,\sigma_{j}^{2}\mathbf{I}_{2})$, with $\sigma_{j}=[0.01,0.1,0.5,1.0,2.0,4.0]$.
3D GRU 2-bit flip-flop (Fig. 5): $\mathcal{N}(\boldsymbol{\mu},\,\sigma_{j}^{2}\mathbf{I}_{3})$, with $\sigma_{j}=[0.01,0.05,0.2,1.0,5.0]$. The mean $\boldsymbol{\mu}$ is chosen as a point on the separatrix found by interpolating between two attractors and performing iterative binary search with ODE (1). A second $\boldsymbol{\mu}$ is obtained by interpolating a different attractor pair for training the second KEF.
11D ecological dynamics (Fig. 6): We similarly identify a single separatrix point $\boldsymbol{\mu}$. Each coordinate $x[i]$ is sampled independently from a Gamma distribution $x[i]\sim\Gamma(\alpha[i],\beta[i])$, where $\alpha[i]$ and $\beta[i]$ are chosen such that the mode of $x[i]$ equals $\mu[i]$ and the variance equals $\sigma_{j}^{2}$, with $\sigma_{j}=[0.01,0.1,0.3,1.0]$.
[…]
$\Sigma=\sigma_{B}^{2}\mathbf{I}_{N}+(\sigma_{A}^{2}-\sigma_{B}^{2})\,\mathbf{u}\mathbf{u}^{\top},$
with
  • $\mathbf{u}\in\mathbb{R}^{N}$ a unit vector along the attractor axis, i.e. $\mathbf{u}=(\mathbf{a}-\mathbf{b})/\|\mathbf{a}-\mathbf{b}\|$, where $\mathbf{a}$ and $\mathbf{b}$ are the two attractors,
  • $\sigma_{A}>0$ the standard deviation along $\mathbf{u}$,
  • $\sigma_{B}>0$ the standard deviation along all orthogonal directions, and
  • $\mathbf{I}_{N}$ the $N\times N$ identity matrix.
As before $\boldsymbol{\mu}$ is the point on the separatrix along the line joining the attractors.
We set $\sigma_{A}$ to include both attractors, and choose $\sigma_{B}$ as large as possible while avoiding spurious basins. For 300 samples drawn from $\mathcal{N}(\boldsymbol{\mu},\Sigma)$, we evolve the dynamics forward for 3.0 s and estimate the covariance of the resulting approximately forward-invariant distribution, denoted $\tilde{\Sigma}$.
Appendix H Scaling of compute with Dimensionality